iYuko (iyuko):

Simplify -1+4/3-6+4^3

iYuko (iyuko):

Plz help!

OpenStudy (anonymous):

175/3

OpenStudy (anonymous):

175 over 3

iYuko (iyuko):

Where 175 come from?

OpenStudy (anonymous):

the lcd

iYuko (iyuko):

y u need the lcd?

iYuko (iyuko):

I need you to explain not give the answer.

OpenStudy (anonymous):

thats what you do in simplfy

TheSmartOne (thesmartone):

Can you write that with the equation editor?

iYuko (iyuko):

Yes

TheSmartOne (thesmartone):

Please do. Thanks :)

iYuko (iyuko):

\[-1+4\div3-6+4^{3}\]

iYuko (iyuko):

I don't know how to solve

TheSmartOne (thesmartone):

Ok, so do you know PEMDAS? We have to follow that here. :)

iYuko (iyuko):

Yea

iYuko (iyuko):

No ( )

OpenStudy (anonymous):

I think its 1

iYuko (iyuko):

Please explain.

TheSmartOne (thesmartone):

So: Parenthesis Exponents Multiplication and division Addition and Subtraction

iYuko (iyuko):

Yes

TheSmartOne (thesmartone):

We have to follow that from left to right. No parenthesis, so we move to exponents.

iYuko (iyuko):

We have an exponent

TheSmartOne (thesmartone):

What is \(4^3=?\)

iYuko (iyuko):

4^3 = 64?

TheSmartOne (thesmartone):

4^3 = 4 x 4 x 4 = 64 Correct (:

iYuko (iyuko):

Now multiplication

iYuko (iyuko):

We don't have any

TheSmartOne (thesmartone):

So then we move on to division from left to right.

iYuko (iyuko):

4/3=1.33333333333333333

TheSmartOne (thesmartone):

Correct, so now we have currently: \(\sf -1 + 1.3333 - 6 + 64\)

iYuko (iyuko):

Yea now addition?

TheSmartOne (thesmartone):

We can do the addition and the subtraction together. So basically just moving left to right. -1 + 1.3333 = ? And then subtract that by 6, and than add that again by 64

iYuko (iyuko):

0.3333

TheSmartOne (thesmartone):

0.333 - 6 = ?

TheSmartOne (thesmartone):

And whatever you get from doing that ^ then just add 64 to it :P

OpenStudy (anonymous):

never mind I di something wrong